​Kathy Smith​, owner of Flower Petal​, operates a local chain of floral shops. Each shop has its own delivery van. Instead of charging a flat delivery​ fee, Smith wants to set the delivery fee based on the distance driven to deliver the flowers. Smith wants to separate the fixed and variable portions of her van operating costs so that she has a better idea how delivery distance affects these costs. She has the following data from the past seven​ months:

Let's begin by determining the formula that is used to calculate the variable cost​ (slope).

Change in cost

/

Change in volume

=

Variable cost (slope)

Now determine the formula that is used to calculate the fixed cost component.

Total operating cost

-

Total variable cost

=

Fixed cost

Use the​ high-low method to determine Flower Paradise Flower Paradise​'s operating cost equation. ​(Round the variable cost to the nearest cent and the fixed cost to the nearest whole​ dollar.)

y = $

x + $

Use the operating cost equation you determined above to predict van operating costs at a volume of

The operating costs at a volume of 16,500 miles is $

.

Month

Miles Driven

Van Operating Costs

January. . . . . . . . . . . . . . .

16,000

$5,490

February. . . . . . . . . . . . . . .

17,500

$5,700

March. . . . . . . . . . . . . . . . .

14,900

$4,910

April. . . . . . . . . . . . . . . . . .

16,200

$5,340

May. . . . . . . . . . . . . . . . . .

16,900

$5,820

June. . . . . . . . . . . . . . . . . .

15,100

$5,410

July. . . . . . . . . . . . . . . . . . .

14,500

$4,920
Answer #1

Variable cost = change in cost / change in volume

= (5700-4920)/(17500-14500)

= 0.26 per mile

Total operating cost - Total Variable cost = Fixed cost

5700= (17500*0.26) - Fixed cost

Fixed cost = 1150

y = 0.26x +$1150

Operating cost for 16,500 miles

= (0.26*16500) + 1150

= 5440